OpenStudy (girlgoyleh.):

factor completely: 4a^3-16a

Vocaloid (vocaloid):

hint: what's the greatest common factor between 4 and 16?

OpenStudy (girlgoyleh.):

4

Vocaloid (vocaloid):

good! now what's the common factor between a and a^3?

OpenStudy (girlgoyleh.):

a?

Vocaloid (vocaloid):

yes! now when we factor out 4a we get 4a(a^2 - 4) know where to go from here?

OpenStudy (girlgoyleh.):

No

Vocaloid (vocaloid):

hint: when we have a difference of squares we can factor it like so: (A^2 - B^2) = (A+B)(A-B) so (a^2 - 4) = ???

OpenStudy (girlgoyleh.):

it can't be simplified further

Vocaloid (vocaloid):

nope, try again

Vocaloid (vocaloid):



Vocaloid (vocaloid):



OpenStudy (girlgoyleh.):

a-4??

Vocaloid (vocaloid):

nope

Vocaloid (vocaloid):

remember: A^2 - B^2 = (A+B)(A-B) a^2 - 4 = ???

Vocaloid (vocaloid):

hint: what does B equal?

OpenStudy (girlgoyleh.):

4

Vocaloid (vocaloid):

nope.

Vocaloid (vocaloid):

hint: B^2 = 4 so B = ?

OpenStudy (girlgoyleh.):

16

OpenStudy (john_doe02125):

When your done Vocaloid, I have a question to ask you.

Vocaloid (vocaloid):

nope... one more time.

Vocaloid (vocaloid):

B squared is 4 what is B?

OpenStudy (girlgoyleh.):

@John_Doe02125 go away and start your own question please, theyre helping me

OpenStudy (girlgoyleh.):

ohhhhhh, b is 2